The Ksp of Ag2CrO4 is 1.1 * 10^–12 at 298K. The solubility (in mol/L) of Ag2CrO4 in a 0.1M AgNO3 solution is

Grade:11

1 Answers

Agni Sain
7 years ago
Write down the dissociation reaction of Ag2CrO4.Ag2CrO4------>2Ag+ + CrO42-Due to common ion effect,Add the moles of AgNO3 to the solubility const. of Ag+Therefore, net Ag+ ions = 2s+0.1. (Let solubility const. be s)Now,(2s+0.1)^2×s=1.1×10^-12As s s=1.1×10^-10